How can i determine if the followingpairs of lines are parallel, perpendicular or neither: (y=-3/4x-2 ) (6x+8y=-5)
Answer by Theo(13342)

y = -3/4x - 2
-----
6x + 8y = -5
-----
Slope Intercept form of the equation for a straight line is y = mx + b where m is the slope and b is the y intercept.
-----
First equation is already in that form.
-----
In the second equation, solve for y in the equation of:
6x + 8y = - 5
Subtract 6x from both sides of this equation to get:
8y = -6x - 5
Divide both sides of this equation to get:
y = (-6x - 5)/8 which becomes:
y = (-6/8)*x - (5/8)
-----
Slope of the first line is -(3/4)
Slope of the second line is -(6/8)
-(3/4) is the same as -(6/8) so the lines are parallel.
-----
Graph of both lines is shown below:

-----
The y intercept of your first line is -2.
The y intercept of your second line is -(5/8)
-----
Slope Intercept form of the equation for your first line is:
y = (-3/4)*x - 2 where (-3/4) is the slope and -2 is the y intercept.
-----
Slope Intercept form of the equation for your second line is:
y = (-6/8)*x - (5/8) where (-6/8) is the slope and -(5/8) is the y intercept.
-----
The equation of the second line can be simplified further to equal:
y = (-3/4)*x - (5/8) where (-3/4) is the slope and -(5/8) is the y intercept.
-----